These were accepted.
Agenda for this month's meeting was full. Will try again at next meeting. Deferred until May.
More progress. Deadline is now end of March.
Deferred until April since Alastair and Stephen are away.
All Units now managing their sites. Alastair to bring to notice of CCPAG by email. Handling of this matter devolved to CEG.
Undeferred. Chris to try out fs flushall on openAFS 1.6.6
This will be rolled into the general AT decant
Blog articles suggested:
Blog articles completed:
Report from Computing Executive Group
There was a general view that it would be desirable to obtain independent confirmation of the correctness of the Gas suppression system (including the provision of a hold-off button) before the system was re-enabled. It was pointed out that the self-managed server room was also covered by this system.
Ian asked about the apparent lack of progress with this and the fact that considerations about further academic involvement appeared to have cropped up out of nowhere. Alastair explained that the slow progress was due to the change in the Head of School and that the academic considerations came out of the deliberations on the Computing Staff's raison d'etre
This was delayed until the next meeting
Reports from units
Last week's failure has left us without any cold spare switch replacements. Suggested Techs be given shot at fixing failed switches. In theory the AT fabric is expendable and could be used as a source of spares should another failure occur. Buying reconditioned kit is also a possibility.
Topics for discussion
It was agreed that the best solution was to provide people with their own VM for this with onward access to the Informatics network disabled. If provisioning of VMs becomes a problem, we may need to revisit this.
Stephen reminded everyone about the ITPF Linux Day on 14th March.
The next meeting will be on 26th March 2014 at 10:00 in IF-4.31, chaired by George.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh